OpenStudy (anonymous):

HELP PLZ! Which fraction is equivalent to 9/12

OpenStudy (sleepyjess):

To get equivalent fractions, multiply the numerator and denominator by the same number. For example, \(\dfrac{9}{12}*\dfrac{3}{3}=\dfrac{27}{36}\). You can also divide to get a smaller fraction, for example \(\dfrac{2}{4}/2=\dfrac{1}{2}\)

OpenStudy (anonymous):

Thx @sleepyjess !! but in my case, what will the number be that i am mutiplying with?

OpenStudy (anonymous):

You can multiply the numerator and the denominator by any number (it has to be the same numbers). Once this is done, you'll get an Equivalent fraction.

OpenStudy (anonymous):

I still dont understand!

OpenStudy (sleepyjess):

Do you have answer choices?